OpenStudy (anonymous):

Help! What is the expression in factored form? x^2 + 14x + 48 A. (x + 6)(x - 8) B. (x + 8)(x - 6) C. (x - 8)(x - 6) D. (x + 6)(x + 8)

hartnn (hartnn):

since all the signs in the question are + all the signs in the answer must also be + :P

OpenStudy (anonymous):

x = -6, -8

OpenStudy (campbell_st):

find the factors of 48 that add to 14 so as an example you can have 1 x 48 = 48...... 1 + 48 = 49 so discard those 2 x 24 = 48 2 + 24 = 26 discard

OpenStudy (anonymous):

C. (x - 8)(x - 6)